OpenStudy (anonymous):

-3|2-5/4u|(less than or equal to sign) -18

OpenStudy (anonymous):

yes .

hero (hero):

Okay

hero (hero):

You can continue solving for u from there

hero (hero):

Is that what the back of your book says?

OpenStudy (anonymous):

It says u <= -3 1/5 or u >= 6 2/5

hero (hero):

Maybe you interpreted the problem wrong. can you scan it directly from the book? Isn't it in your Algebra book? I still have that. Tell me the section and page number.

OpenStudy (anonymous):

Page 421, problem 17 .

hero (hero):

Something is very weird about it because -3 1/5 = -16/5 and 6 2/5 = 32/5 It's like backwards or something

OpenStudy (anonymous):

Maybe you can check it out .

hero (hero):

I will

hero (hero):

Chrissy, give me the exact page number because I'm having trouble finding it. Give me the actual book page number not the adobe page number, thanks

hero (hero):

Remember, I'm using adobe, so..., but yeah, I can't find it, however, I'm in the right section, I just need you to double-check the page number

hero (hero):

And also, tell me the name of the section of the book.

hero (hero):

I will fix the solution

OpenStudy (anonymous):

Okay .

hero (hero):

\[-3|2-\frac{5u}{4}| \le -18\]

hero (hero):

:( I was going to fix the solution myininaya. It was an honest mistake where I put u in the denominator

myininaya (myininaya):

\[|2-\frac{5u}{4}| \ge 6\] \[|8-5u| \ge 24 \text{ note I multiplied both sides by 4}\] \[|f(x)| \ge 24 => f(x) \ge 24 \text{ or } f(x) \le -24\] So we have \[8-5u \ge 24 \text{ or } 8-5u \le -24\] \[-5u \ge 16 \text{ or } -5u \le -32\] \[u \le \frac{-16}{5} \text{ or } u \ge \frac{32}{5}\]

hero (hero):

I wish you had let me fix it

myininaya (myininaya):

You can still do it. You want me delete this?

hero (hero):

No, it's too late now.

OpenStudy (anonymous):

I understand it . Thanks guys . :)